OpenStudy (anonymous):

Consider the functions f(x) = x2 − 13 and g(x) = x + 5. What is the value of f[g(−4)]? −12 3 −13 8

OpenStudy (solomonzelman):

\(\LARGE\color{orangered}{ f(x) = x^2 − 13 }\) \(\LARGE\color{orangered}{ g(x) = x + 5 }\) \(\LARGE\color{orangered}{ f(g(x))= (x+5)^2 − 13 }\) \(\LARGE\color{orangered}{ f(g(x))=x^2+10x+25 − 13 }\) \(\LARGE\color{orangered}{ f(g(x))=x^2+10x+12 }\) \(\LARGE\color{orangered}{ f(g(-4))=(-4)^2+10(-4)+12 }\) \(\LARGE\color{orangered}{ f(g(-4))=16-40+12 }\) \(\LARGE\color{orangered}{ f(g(-4))=-14 }\)